Find the interest rate needed for an investment of $5600 to double in 7 yr if interest is compounded continuously. (Round your answer to two decimal places.)

Answer by lwsshak3(6469)

Find the interest rate needed for an investment of $5600 to double in 7 yr if interest is compounded continuously. (Round your answer to two decimal places.)
**
Compound interest formula for continuous compounding: A=Pe^rt, P=initial amt, r=annual interest rate, A=amt after t years.
For given problem:
A=Pe^rt
A/P=e^rt=2
e^7r=2
take log of both sides
7r*lne=ln2
lne=1
r=ln2/7
r≈9.90%
note: initial investment of $5600 not needed for solution